# When Robustness Meets Conservativeness: Conformalized Uncertainty Calibration for Balanced Decision Making

### v1

@@ 3.2 Theoretical Analysis
We establish two key theoretical properties of the proposed estimator. First, we prove its validity, showing that under the exchangeability assumption, the proposed estimator $\hat{\alpha}_{\ell}(\lambda)$ upper bounds the true risk for any specification of the robustness parameter $\lambda$. Then, we demonstrate that under a majorant consistency assumption, the conservative trajectory traced by the estimator $(\hat{\alpha}_{I}(\lambda),\hat{\alpha}_{R}(\lambda))$ by varying $\lambda$ characterizes a certified Pareto frontier.
Theorem 1 (Validity).
Under 1 and 2, the estimator $\hat{\alpha}_{\ell}(\lambda)$ in (5) satisfies
$\mathbb{E}[\hat{\alpha}_{\ell}(\lambda)]\geq\mathbb{E}[\ell_{\lambda}(X_{n+1},Y_{n+1})].$
Proof of Theorem 1.
We prove by showing:
$\mathbb{E}\left[\hat{\alpha}_{\ell}(\lambda)\right]\overset{(a)}{\geq}\mathbb{E}\left[\tilde{\alpha}_{\ell}(\lambda)\right]\overset{(b)}{\geq}\mathbb{E}[\ell_{\lambda}(X_{n+1},Y_{n+1})].$
[…]
therefore $\hat{\alpha}_{\ell}(\lambda)\geq\tilde{\alpha}_{\ell}(\lambda)$ almost surely, which implies $(a)$. To prove $(b)$, we expand the difference of two sides of $(b)$ by using the definition of $\tilde{\alpha}_{\ell}(\lambda)$ in (6):
$~\mathbb{E}\left[\tilde{\alpha}_{\ell}(\lambda)\right]-\mathbb{E}[\ell_{\lambda}(X_{n+1},Y_{n+1})]$
$=$ $~\left(\frac{n}{n+1}-1\right)\mathbb{E}[\ell_{\lambda}(X_{n+1},Y_{n+1})]+\frac{B}{n+1}$
$=$ $~\frac{B-\mathbb{E}[\ell_{\lambda}(X_{n+1},Y_{n+1})]}{n+1}\geq 0.$
where the first equation follows from 1, and the inequality follows from 2. This proves $(b)$ holds and thus concludes the proof. ∎
We note that our notion of validity resembles the idea of post-hoc validity studied in the literature, such as [Gauthier et al., 2025a]. In both cases, the prediction sets are specified first, and the analysis then determines the corresponding risk guarantee $\alpha$. However, there is a key difference: In prior work, the candidate set is typically data-dependent, and validity is ensured through post-hoc adjustment, often relying on e-value methods [Balinsky and Balinsky, 2024, Gauthier et al., 2025a, Gauthier et al., 2025b]. By contrast, in our framework the robustness level $\lambda$—and hence the uncertainty set $\mathcal{C}_{\lambda}$—is specified a priori, independent of the data. The estimator $\hat{\alpha}_{\ell}(\lambda)$ then directly provides a guaranteed bound on the associated risk without adaptive tuning. As a result, our method preserves validity with substantially higher accuracy, i.e., a smaller deviation from the true risk, compared to e-value–based approaches.
Additionally, under a slightly more demanding i.i.d. conditioning, we can also establish a finite-sample error bound for the proposed estimator.
Proposition 2 (Finite-sample error bound).
Under 2 and suppose $\{(x_{i},y_{i})\}_{i=1}^{n+1}$ are i.i.d. sampled, then for any $\delta>0$ and $\lambda\in\Lambda$, with probability at least $1-\delta$,
@@
Proposition 2 is useful, as when combined with Theorem 1, we can now a valid one-shot estimator with at least probability $1-\delta$ by offsetting it with the error defined in (7):
@@
Next, we show that the trajectory describing the miscoverage-regret trade-off is (weakly) decreasing under the standard majorant consistency assumption [Delage and Ye, 2010, Rahimian and Mehrotra, 2019, Mohajerin Esfahani and Kuhn, 2018]. This assumption is crucial for guaranteeing that the trade-off curve preserves a monotone structure, thereby ensuring the existence of a well-defined Pareto frontier.
Assumption 3 (Majorant consistency).
@@
3 describes the condition that shrinking the uncertainty set does not increase the conditional expected realized cost. It holds in a broad range of standard optimization problems, including robust linear optimization with polyhedral or ellipsoidal uncertainty sets [Ben-Tal and Nemirovski, 1999], robust quadratic optimization, regression with convex loss functions (e.g., squared, absolute, or Huber loss) under norm-ball uncertainty [Ben-Tal and Nemirovski, 2001], and robust classification with convex surrogate losses (e.g., hinge or logistic) [Bertsimas et al., 2011]. It is also directly implied by distributionally robust optimization (DRO) formulations with Wasserstein or $\phi$-divergence balls, where monotonicity of the robust objective in the size of the ball is well established [Delage and Ye, 2010, Rahimian and Mehrotra, 2019, Mohajerin Esfahani and Kuhn, 2018]. We present an example and a counterexample for 3 in Appendix A.
Combined with the properties of nested uncertainty sets and bounded regret in our problem setting, we can prove that the image of the trade-off curve coincides with the Pareto frontier of the robust policy family. The proof can be found in Appendix C.
Proposition 3 (True Pareto frontier).
Let $\alpha_{I}(\lambda)\coloneqq\mathbb{E}[I_{\lambda}(X,Y)]$ and $\alpha_{R}(\lambda)\coloneqq\mathbb{E}[R_{\lambda}(X,Y)]$. Under Assumption 3, for any $\lambda_{1}\leq\lambda_{2}$,
@@
Corollary 1 (Certified Pareto Frontier).
@@
By Corollary 1, the lower-left envelope of $\widehat{\mathcal{F}}$ traces the certified Pareto frontier, thus providing guaranteed trade-offs between regret and miscoverage.
@@ 4 Experiments
Figure 4: Miscoverage–regret tradeoff Pareto frontiers with attained solutions from each model. For a prespecified preference weighting, the black star denotes the true optimal tradeoff. The red dot indicates the tradeoff obtained from the estimated $\hat{\lambda}$ using CREME. Blue markers correspond to three baseline methods that identify $\lambda$ values ensuring the miscoverage rate remains below $5\%$ (vertical dashed blue line).
This section evaluates the performance of our proposed algorithm in Algorithm 1, termed Conformal REgret Miscoverage Estimate (CREME). We focus on four representative optimization paradigms: ($i$) linear programming; ($ii$) the newsvendor problem (single-item inventory); (iiiiii) portfolio optimization (long-only, mean–CVaR proxy); and ($iv$) the shortest path problem (equivalently, a unit flow problem). These problems are chosen because they exemplify widely studied optimization settings, and their regret can be efficiently computed using convex solvers. In each case, we designate certain parameters in the optimization objective (e.g., the cost vector) as the uncertain random variable $Y$, and the uncertainty set is defined as a $\ell_{\infty}$ ball centered at $\mathbb{E}[Y]$. Additional information for experimental configurations and omitted details in the remaining subsections are all deferred Appendix D.
@@ 4.1 Validity-Accuracy Analysis
We first evaluate the quality of the miscoverage-regret frontier produced by CREME using two metrics: validity and accuracy. Validity [Lei et al., 2018] measures whether the estimator successfully upper bounds the true risk, while accuracy assesses the tightness of this bound. Formally, validity is defined as the proportion of trials in which both $\hat{\alpha}_{R}(\lambda)$ and $\hat{\alpha}_{I}(\lambda)$ satisfy their validity conditions. Accuracy [Zhou et al., 2025] is defined as the negative average Euclidean distance between the estimated front $(\hat{\alpha}_{I}(\lambda),\hat{\alpha}_{R}(\lambda))$ and the corresponding ground-truth front $(\alpha_{I}(\lambda),\alpha_{R}(\lambda))$, i.e., negative of the mean error. When calculating these metrics, the ground truth front is approximated via Monte Carlo sampling of size $100$ from the true distribution and averaged over $20$ independent trials. Throughout our evaluation, $\lambda$ is varied from zero to the support radius of $Y$.
@@
Figure 3 presents the evaluation results, comparing CREME with the empirical Monte Carlo estimator (i.e., $\bar{\ell}_{n}(\lambda)$). It can be seen that across all optimization settings, the Monte Carlo estimator (gray) tends to lie toward the left of the plot, while CREME (red) concentrates in the upper-right region. This difference arises because naive baselines are primarily designed for consistent estimation, achieving high accuracy without validity guarantees. In contrast, our proposed procedure strikes a balanced trade-off, maintaining strong performance on both validity and accuracy, and tracing a Pareto frontier [^1] This should be distinguished from the miscoverage-regret Pareto frontier in this paper. over these two dimensions. These results not only support our theoretical analysis but also highlight that CREME is particularly well-suited for high-stakes decision-making tasks where both accuracy and validity are critical.
@@
In particular, Figure 3 also provides deeper insights into the properties of CREME associated with the number of calibration data, which is shown in its performance trajectory (red dashed line). In the small-sample regime where there is limited calibration data, CREME exhibits relatively low accuracy but high validity, positioning itself in the lower-right region of the plot. This behavior is desirable: under uncertainty and limited resources, CREME prioritizes conservatism by erring on the side of validity. As the calibration sample size increases, the trajectory rises sharply along the accuracy axis (at an almost exponential rate) while declining only gradually along the validity axis (at an approximately linear rate). This imbalance in rates is crucial: it ensures the existence of a Pareto frontier and enables CREME to maintain favorable trade-offs between validity and accuracy. In the large-sample regime, the trajectory converges toward the region occupied by the Monte Carlo estimator. This is expected, as our method is by design asymptotically consistent with the Monte Carlo estimator, explaining both the gradual deterioration of validity and the simultaneous improvement in accuracy as calibration size increases. Together, these observations provide a clear view into the internal behavior of CREME, demonstrating how it adapts across regimes and achieves strong trade-offs between accuracy and validity.
@@ 4.2 Decision Quality Evaluation
We also assess the quality of the robustness parameter $\hat{\lambda}$ selected by CREME. We show that given any pre-specified weight that encodes the decision-maker’s preference for the miscoverage–regret tradeoff, CREME can identify $\hat{\lambda}$ that closely approximates the optimal $\lambda^{*}$, thereby achieving tradeoff $)(\hat{\alpha}_{I}(\hat{\lambda}),\hat{\alpha}_{R}(\hat{\lambda}$)) that approximates the optimal tradeoff $(\hat{\alpha}_{I}(\lambda^{*}),\hat{\alpha}_{R}(\lambda^{*}))$.
@@
Figure 4 presents the evaluation results. It can be seen that the estimated Pareto frontier (red) aligns closely with the true frontier (black), lying only slightly above it. This provides a clean visual supporting evidence to Figure 3 that our method is both valid and accurate. Furthermore, observe that in the newsvendor and portfolio selection problems, the ground-truth curves traced from varying $\lambda$ contain points that are not strictly on the Pareto frontier (i.e., non-dominant points), whereas the linear programming and shortest path problems yield well-defined Pareto frontiers. This observation is consistent with 3 and Proposition 3, where we argued that only some particular classes (e.g., linear) are guaranteed to have full Pareto frontiers.
Next, observe from Figure 4 that the tradeoff points selected by the three baseline methods deviate substantially—achieving low miscoverage but suffering from high regret. In contrast, the point selected by CREME (red dot) lies very close to the true optimal tradeoff point (black star), striking a balanced compromise between miscoverage and regret. This improvement stems from two factors. First, unlike the baselines that are designed to conservatively control miscoverage and thus myopically select suboptimal robustness parameters, CREME explicitly accounts for both miscoverage and regret, enabling it to identify a globally balanced solution. Second, the high accuracy of our estimator (as previously discussed) allows CREME to closely trace the true Pareto frontier, thereby preserving the correct relative positioning of tradeoff points and guiding the selection of appropriate robustness parameters. Together, these results demonstrate that CREME effectively supports decision-makers in selecting robustness levels that achieve a desired balance between miscoverage and regret.

### v2

@@ 3.2 Theoretical Analysis
We establish two key theoretical properties of the proposed estimator. First, we prove its validity, showing that under the exchangeability assumption, the proposed estimator $\hat{\alpha}_{\ell}(\lambda)$ upper bounds the true risk for any specification of the robustness parameter $\lambda$. Then, we demonstrate that under a majorant consistency assumption, the conservative trajectory traced by the estimator $(\hat{\alpha}_{I}(\lambda),\hat{\alpha}_{R}(\lambda))$ by varying $\lambda$ characterizes a certified Pareto frontier. All proofs in this section can be found in Appendix A through Appendix D.
Theorem 3.4 (Validity).
Under 3.1 and 3.2, the estimator $\hat{\alpha}_{\ell}(\lambda)$ in (5) satisfies
$\mathbb{E}[\hat{\alpha}_{\ell}(\lambda)]\geq\mathbb{E}[\ell_{\lambda}(X_{n+1},Y_{n+1})].$ (7)
Additionally, under the i.i.d. condition, we can establish a finite-sample error bound for the proposed estimator.
Proposition 3.5 (Finite-sample error bound).
Under 3.2 and suppose $\{(x_{i},y_{i})\}_{i=1}^{n+1}$ are i.i.d., then for any $\delta>0$ and $\lambda\in\Lambda$, with probability at least $1-\delta$,
@@
Proposition 3.5 can be combined with Theorem 3.4 to produce a valid one-shot estimator with high probability by offsetting it with the error $\epsilon$ defined in (8):
@@
Next, we show that the miscoverage-regret trade-off curve is (weakly) decreasing under the standard majorant consistency assumption (Delage and Ye, 2010; Rahimian and Mehrotra, 2019; Mohajerin Esfahani and Kuhn, 2018). This assumption is crucial for guaranteeing that the trade-off curve preserves a monotone structure, thereby ensuring the existence of a well-defined Pareto frontier.
Assumption 3.6 (Majorant consistency).
@@
3.6 describes the condition that shrinking the uncertainty set does not increase the conditional expected realized cost. It holds in a broad range of standard optimization problems, including robust linear optimization with polyhedral or ellipsoidal uncertainty sets (Ben-Tal and Nemirovski, 1999), robust quadratic optimization, regression with convex loss functions (e.g., squared, absolute, or Huber loss) under norm-ball uncertainty (Ben-Tal and Nemirovski, 2001), and robust classification with convex surrogate losses (e.g., hinge or logistic) (Bertsimas et al., 2011). It is also directly implied by DRO formulations with Wasserstein or $\phi$-divergence balls, where monotonicity of the robust objective in the size of the ball is well established (Delage and Ye, 2010). For more intuition, we provide an example and a counterexample of 3.6 in Appendix C.
Combined with the properties of nested uncertainty sets and bounded regret in our problem setting, we can prove that the image of the trade-off curve coincides with the Pareto frontier of the robust policy family.
Proposition 3.7 (True Pareto frontier).
Let $\alpha_{I}(\lambda)\coloneqq\mathbb{E}[I_{\lambda}(X,Y)]$ and $\alpha_{R}(\lambda)\coloneqq\mathbb{E}[R_{\lambda}(X,Y)]$. Under Assumption 3.6, for any $\lambda_{1}\leq\lambda_{2}$,
@@
Corollary 3.8 (Certified Pareto frontier).
@@
By Corollary 3.8, the lower-left envelope of $\widehat{\mathcal{F}}$ traces the certified Pareto frontier, thus providing guaranteed trade-offs between regret and miscoverage.
3.3 Re-calibration After Selection
Algorithm 1 CREME
[…]
3:for $j\in\left\{1,2\right\}$ do
4:  for $\lambda\in\{\lambda_{i}\}$ do
5:   $\bar{I}_{n}^{(j)}(\lambda)\leftarrow\sum_{i\in\mathcal{I}_{i}}I_{\lambda}(x_{i},y_{i})/n$;
6:   $\bar{R}_{n}^{(j)}(\lambda)\leftarrow\sum_{i\in\mathcal{I}_{i}}R_{\lambda}(x_{i},y_{i})/n$;
7:   $\tilde{\alpha}_{R}^{(j)}(\lambda)\leftarrow$ compute (6) by $\bar{R}_{n}^{(j)}(\lambda)$, $B=R_{\text{max}}$;
8:   $\tilde{\alpha}_{I}^{(j)}(\lambda)\leftarrow$ compute (6) by $\bar{I}_{n}^{(j)}(\lambda)$, $B=1$;
9:   $\hat{\mathcal{F}}^{(j)}\leftarrow\hat{\mathcal{F}}^{(j)}\cup\{(\hat{\alpha}_{I}^{(j)}(\lambda),\hat{\alpha}_{R}^{(j)}(\lambda))\}$;
10:  end for
11:  Remove dominated points from $\hat{\mathcal{F}}^{(j)}$;
12:end for
[…]
14:return Post-hoc valid estimate $)(\hat{\alpha}_{I}^{(2)}(\hat{\lambda}),\hat{\alpha}_{R}^{(2)}(\hat{\lambda}$)).
In the previous sections, our analysis was carried out under the assumption that the robustness parameter $\lambda\in\Lambda$ is specified a priori. Under this setting, our estimator is used purely for risk assessment: for each candidate robustness level $\lambda$, it provides conservative estimates of the miscoverage and regret that would be incurred by deploying the corresponding robust decision. Our theoretical results establish that, when $\lambda$ is fixed in advance, the estimator enjoys exact finite-sample validity guarantees.
In practical decision-making settings, however, it is often desirable for the robustness parameter to be chosen a posteriori, after inspecting the estimated trade-offs. For example, after observing the estimated Pareto frontier $\hat{\mathcal{F}}$, a decision-maker may select a robustness level that optimizes a subjective miscoverage–regret preference,
$\hat{\lambda}=\operatorname*{arg~max}_{\lambda\in\Lambda}g\left(\hat{\alpha}_{I}(\lambda),\hat{\alpha}_{R}(\lambda)\right),$ (10)
where $g:\mathbb{R}^{2}\to\mathbb{R}$ encodes the decision-maker’s risk preference. The selected robustness level $\hat{\lambda}$ is then deployed, and the corresponding estimated risks are reported.
This post-hoc selection step fundamentally alters the statistical setting. Because $\hat{\lambda}$ is chosen based on the same calibration data used to estimate risk, the loss sequence $\{\ell_{\hat{\lambda}}(X_{i},Y_{i})\}_{i=1}^{n+1}$ is no longer exchangeable. As a result, the exact validity guarantee of Theorem 3.4 may fail to hold. The following corollary formalizes this loss of validity.
Corollary 3.9 (Post-hoc validity degradation).
[…]
$\mathbb{E}\bigl[\hat{\alpha}_{\ell}(\hat{\lambda})\bigr]\geq\mathbb{E}\bigl[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})\bigr]-\Delta(g,\mathcal{P}),$
where the error term is defined as
$\Delta(g,\mathcal{P})\coloneqq\sup_{1\leq i\leq n}\biggl\lvert\mathbb{E}\bigl[\ell_{\hat{\lambda}}(X_{i},Y_{i})\bigr]-\mathbb{E}\bigl[\ell_{\hat{\lambda}}(X_{n+1},Y_{n+1})\bigr]\biggr\rvert$.
Intuitively, the error term $\Delta(g,\mathcal{P})$ depends on how much asymmetry that selecting $\hat{\lambda}$ by $g$ introduces to the distribution $\mathcal{P}$. It arises because the data-dependent choice of $\hat{\lambda}$ leaks information from the calibration sample into the loss evaluation, thereby breaking the symmetry between calibration points and the test point. This violation of exchangeability is well known to undermine conformal-style guarantees (Barber et al., 2023). Consequently, the estimator may no longer be conservative after post-hoc selection. A detailed proof is provided in Appendix E.
We address this issue using a simple and effective data-splitting strategy (Cox, 1975; Rasines and Young, 2023), which restores exact finite-sample validity under post-hoc selection. Specifically, we randomly partition the calibration dataset $\{(X_{i},Y_{i})\}_{i=1}^{n}$ into two disjoint subsets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. Using $\mathcal{D}_{1}$, we construct a pre-hoc estimator $\hat{\alpha}_{\ell}^{(1)}(\lambda)$ and form the corresponding Pareto frontier $\hat{\mathcal{F}}^{(1)}$. Only this pre-hoc Pareto frontier is revealed to the decision-maker and used to select the robustness level $\hat{\lambda}$ via (10). Independently, using $\mathcal{D}_{2}$, we construct a post-hoc estimator $\hat{\alpha}_{\ell}^{(2)}(\lambda)$, which is then evaluated at the selected robustness level $\hat{\lambda}$ to produce the re-calibrated miscoverage and regret estimates $)\big(\hat{\alpha}_{I}^{(2)}(\hat{\lambda}),\hat{\alpha}_{R}^{(2)}(\hat{\lambda}$$\big)$. The pseudocode in Algorithm 1 summarizes our entire proposed method, which we term as Conformal REgret Miscoverage Estimate (CREME).
@@ 4 Experiments
Figure 4: Miscoverage-regret tradeoff Pareto frontiers. Given some prespecified preference function, there is an optimal tradeoff point (black star) that can be computed from the true Pareto frontier (black line). We can obtain a certified tradeoff point (red dot) and a post-hoc tradeoff point (green point) by using CREME. Blue markers represent tradeoff points derived from the three baseline methods that identify $\lambda$ values ensuring the miscoverage rate remains below $5\%$ (vertical dashed blue line).
This section presents the numerical evaluation of CREME (Algorithm 1). We focus on four representative optimization paradigms for decision-making problems: ($i$) linear programming; ($ii$) the newsvendor problem; (iiiiii) portfolio optimization; and ($iv$) the shortest path problem. These problems are chosen because they exemplify widely studied optimization settings, and their regret can be efficiently computed using convex solvers. Their supremum $B$ defined in 3.2 is either known by design or approximated through the empirical maximum $\max_{1\leq i\leq n}\ell_{\lambda}(x_{i},y_{i})$. Additional information for experimental configurations and details omitted in this section can be found in Appendix F.
@@ 4.1 Validity-Accuracy Analysis
We first evaluate the quality of the miscoverage-regret frontier produced by CREME using two metrics: validity and accuracy. Validity measures whether the estimator successfully upper bounds the true risk, defined as the proportion of trials in which both $\hat{\alpha}_{R}(\lambda)$ and $\hat{\alpha}_{I}(\lambda)$ satisfies (7). Accuracy assesses the tightness of this bound, defined as the negative average distance between $(\hat{\alpha}_{I}(\lambda),\hat{\alpha}_{R}(\lambda))$ and $(\alpha_{I}(\lambda),\alpha_{R}(\lambda))$. They are computed from averaging across $20$ independent trials.
@@
Figure 3 presents the evaluation results comparing CREME with the empirical Monte Carlo estimator (i.e., $\bar{\ell}_{n}(\lambda)$). It can be seen that across all optimization settings, the Monte Carlo estimator (gray) tends to lie toward the left of the plot, while CREME (red and green) concentrates in the upper-right region. The post-hoc certified estimation (green) consistently remains at a higher validity compared to its standard counterpart (red), but with a slightly degraded accuracy. This difference arises because naive baselines are primarily designed for consistent estimation, achieving high accuracy without validity guarantees. In contrast, our proposed procedure strikes a balanced trade-off, maintaining strong performance on both validity and accuracy. These results highlight that CREME is well-suited for high-stakes decision-making tasks where both accuracy and validity are critical.
@@
Figure 3 also provides insights into the finite-sample behavior of CREME. In the small-sample regime with a small $n$, CREME exhibits relatively low accuracy but high validity, positioning itself in the lower-right region of the plot. This behavior is desirable: under uncertainty and limited resources, CREME prioritizes conservatism by erring on the side of validity. As $n$ increases, the trajectory rises at around an exponential rate along the accuracy axis while declining only at roughly a linear rate along the validity axis. This difference in rates ensures the existence of a Pareto frontier and enables CREME to maintain favorable trade-offs between validity and accuracy. In the large-sample regime, the trajectory converges toward the region occupied by the Monte Carlo estimator, demonstraing asymptotic consistency.
Table 1: Sensitivity analysis of CREME against number of samples $n$ and the robustness level index set size $|\Lambda|$.
Linear Programming Newsvendor Portfolio Optimization Shortest Path Gap Time (s) Gap Time (s) Gap Time (s) Gap Time (s) $|\Lambda|=10$ $n=10$ $0.10\pm 0.03$ $0.58\pm 0.03$ $19\pm 0.08 0.0$$00\pm 0.00$ $0.16\pm 0.07$ $0.84\pm 0.10$ $0.11\pm 0.03$ $0.72\pm 0.04$ $n=20$ $0.09\pm 0.03$ $0.81\pm 0.03$ $12\pm 0.04 0.0$$00\pm 0.00$ $0.13\pm 0.07$ $1.14\pm 0.12$ $0.06\pm 0.01$ $0.99\pm 0.06$ $n=30$ $0.08\pm 0.01$ $1.00\pm 0.04$ $13\pm 0.04 0.0$$0\pm 0.00 0.0$$09\pm 0.02$ $1.50\pm 0.13$ $0.06\pm 0.01$ $1.20\pm 0.07$ $|\Lambda|=20$ $n=10$ $0.10\pm 0.03$ $0.58\pm 0.03$ $19\pm 0.08 0.0$$00\pm 0.00$ $0.16\pm 0.07$ $0.84\pm 0.10$ $0.11\pm 0.03$ $0.72\pm 0.04$ $n=20$ $0.09\pm 0.03$ $0.81\pm 0.03$ $12\pm 0.04 0.0$$00\pm 0.00$ $0.13\pm 0.07$ $1.14\pm 0.12$ $0.06\pm 0.01$ $0.99\pm 0.06$ $n=30$ $0.08\pm 0.01$ $1.00\pm 0.04$ $13\pm 0.04 0.0$$0\pm 0.00 0.0$$09\pm 0.02$ $1.50\pm 0.13$ $0.06\pm 0.01$ $1.20\pm 0.07$ $|\Lambda|=30$ $n=10$ $0.09\pm 0.03$ $1.73\pm 0.06$ $16\pm 0.07 0.0$$00\pm 0.00$ $0.17\pm 0.08$ $2.50\pm 0.28$ $0.11\pm 0.04$ $2.13\pm 0.08$ $n=20$ $0.08\pm 0.02$ $2.42\pm 0.09$ $11\pm 0.04 0.0$$00\pm 0.00$ $0.13\pm 0.07$ $3.43\pm 0.34$ $0.05\pm 0.01$ $2.93\pm 0.15$ $n=30$ $0.06\pm 0.02$ $2.98\pm 0.13$ $11\pm 0.04 0.0$$0\pm 0.00 0.0$$08\pm 0.03$ $4.49\pm 0.44$ $0.04\pm 0.01$ $3.63\pm 0.25$
@@ 4.2 Decision Quality Evaluation
We also assess the quality of the robustness parameter selected based on CREME. We show that given any pre-specified weight that encodes the decision-maker’s preference for the miscoverage-regret tradeoff, CREME can identify $\hat{\lambda}$ that closely approximates the optimal $\lambda^{*}$, providing high-quality robustness calibration for decision-making.
@@
Figure 4 visualizes the evaluated Pareto frontier and tradeoff outcome. It can be seen that the estimated Pareto frontiers (red and green) lie closely above the true frontier (black), with the post-hoc certified frontier (green) consistently lying slightly higher above its standard counterpart, providing a slightly conservative but safer miscoverage-regret tradeoff characterization. Both selects robustness level that is close to $\lambda^{*}$ (black dot) and showcase similar miscoverage-regret tradeoff. On the other hand, the tradeoff points selected by the three baseline achieves low miscoverage but suffer from high regret. This is because these baselines, similar to many existing robust set specifying heuristics, are designed to conservatively control miscoverage and thus myopically select suboptimal robustness parameters, leading to an undesirable miscoverage-regret tradeoff selection. On the other hand, CREME is capable of guiding the selection of appropriate robustness parameters that closely resemble its true optimal value, supporting a balanced decision calibration.
On a side note, observe in Figure 4 that the ground-truth curve for the newsvendor and portfolio selection problems contains points that are not strictly on the Pareto frontier (i.e., non-dominant points). This observation consistent with 3.6 and Proposition 3.7, so additional pruning must be done if one desires the full Pareto frontier.
4.3 Ablation Studies
Finally, we conduct an ablation study on CREME’s performances under different hyperparameter specifications: the number of samples $n$ and the robustness level index set size $|\Lambda|$. We examine two performance metrics: ($i$) Gap: the approximation gap of the estimated Pareto frontier to its true counterpart, measured by their average distance (i.e., negative of the accuracy metric); ($ii$) Time (s): the execution time (seconds) for CREME to construct a full estimated Pareto frontier. Numbers are reported as mean $\pm$ standard deviation computed from $20$ trials.
Table 1 summarizes the experimental results. As $n$ and $|\Lambda|$ increase, the approximation gap decreases, consistent with our theoretical bound in Proposition 3.5. At the same time, larger $n$ and $|\Lambda|$ lead to increased runtime, which peaks at approximately $5$ seconds across all configurations. This computational cost is typically manageable in practice since the construction of the frontier is usually a one-shot procedure. In particular, the runtime for the newsvendor setting remains uniformly low, owing to the availability of closed-form solutions. Overall, these results demonstrate that our approach achieves both strong empirical performance and computational efficiency suitable for practical deployment.